# Spatially focused magnetic hyperthermia: comparison of MRSh and sLLG equations

### v1

@@ I Introduction
Magnetic nanoparticle hyperthermia is an adjuvant treatment strategy in cancer therapy where injected magnetic nanoparticles (MNPs) are used to increase the temperature locally, see, e.g. pankhurst ; ortega_d ; pankhurst_progress ; perigo_e_a ; cabrera_d ; clinapp_1 ; clinapp_2 ; review_mnp . The capacity of metallic nanoparticles to preferentially accumulate in tumor tissues is the primary benefit of using them in hyperthermia. This is due to the tumor’s enhanced permeability and retention (EPR) impact, which enables nanoparticles to penetrate and accumulate in the tumor location, for more details and references see, e.g., thesis_mehmeti . The abnormal design of tumor blood vessels, which are leaky and have irregular shapes, causes the EPR effect, enabling nanoparticles to travel through the gaps and collect in the tumor tissue. However, there are several difficulties and limitations linked to magnetic nanoparticle hyperthermia. One of the most difficult tasks is optimizing the nanoparticle characteristics and external energy source to obtain the best therapeutic results. This involves finding the best MNPs size, form, surface coating, and concentration, as well as the best energy dosage and frequency.
@@
Another challenge is to ensure the safety of the procedure, as overheating can harm healthy tissues and result in adverse effects such as discomfort, inflammation, or tissue necrosis. Thus, there is a need for spatially focused heating, which can be done by using the combination of static gradient (DC) and time-dependent alternating (AC) magnetic fields, where the temperature increase is observed only where the static field vanishes mpi_test . The improvement of the efficiency and the spatial focusing ability of MNP hyperthermia receives considerable attention; see, e.g., adv_func_mat ; review_mnp and non_local_heat ; focused_hyperthermia ; recent_focused_hyperthermia ; stat_along_increased . It is known that the transferred energy is given by a bell-shaped curve as a function of the DC field amplitude, see e.g. Fig. (10.4) of review_mnp or superlocal_jmmm ; superlocal_jpd and for more details MRSh_theory_focused .
@@
The energy transferred in a single cycle of the AC field is identical to the area of the dynamic hysteresis loop, and it is related to the imaginary part of the frequency-dependent susceptibility. Measurements of dynamic hysteresis loops under the influence of AC and DC magnetic fields were reported in several publications, see, for example, Mehdaoui ; Myrovali ; Onodera . Similarly, a theoretical study of the dynamics of magnetization in the presence of AC and DC fields was carried out, for example, in Dejardin ; Mrabti ; Murase ; MRSh_theory_focused and recently, the study of the influence of static magnetic field on the specific absorption rate of the randomly oriented assembly of magnetic nanoparticles was reported in Rytov . In addition, the influence of anisotropy and magneto-dipole interaction effects was investigated in the presence of AC and DC fields; see, e.g., usov_claster ; Zhao ; Usov_2019 ; viscous_rotating ; Engelmann .
@@
The single-domain MNP dissipates heat as a result of relaxation losses under the applied AC magnetic field. This heat generation occurs due to either the physical rotation of the particle as a whole (Brownian relaxation) or the internal rotation of the magnetic dipole (Néel relaxation) while the particle is assumed to be fixed brown_neel_relax_1 ; brown_neel_relax_2 ; brown_neel_relax_3 . Although various mathematical approaches are necessary for a theoretical description of the energy absorption processes of immobilized and freely rotating nanoparticles, the usual theoretical framework is the Martsenyuk-Raikher-Shliomis (MRSh) mrsh equation for the Brownian and the stochastic Landau-Lifshitz-Gilbert (sLLG) sLLG equation for the Néel relaxation.
@@
To do that extension, one possible choice is to use the MRSh equation and compare previous results of the sLLG and the MRSh equations directly. Our reasons for this choice are the following.
It is well-known that under some constraints, the two different regimes, i.e. the Brownian and the Néel relaxations are very much separated by each other because the relaxation time $\tau$ is either dominated by $\tau_{{\mathrm{B}}}$ (Brown relaxation time) or by $\tau_{{\mathrm{N}}}$ (Néel relaxation time). So, it is a good approximation to use either the MRSh or the sLLG equations.
However, by using the concept of magnetic and ordinary viscosity, the results of the sLLG equation can, in principle, be used to describe not just the motion of the magnetic dipole but also the particle. Of course, in this case the parameters (damping factor, frequency, magnetic field amplitude etc.) must be chosen appropriately, and these are not necessarily the same if one wants to obtain the same results by the two equations.
Indeed, one of our goals here is to show that one can reproduce well-known dynamic hysteresis loops and bell-shaped spatially-focused energy loss results of the MRSh approach by the sLLG equation. In particular, we demonstrate this by reproducing the MRSh results given on Fig. 2 and Fig. 3 of Ref. MRSh_theory_focused in the framework of the sLLG equation. For example, the spatial focusing ability is different in the right (low frequencies) and left (high frequencies) panels of Fig. 2 of Ref. MRSh_theory_focused . In this work, we show that this is the consequence of the use of a parallel combination of DC and AC magnetic fields and the breakdown of the linear response theory for low frequencies which is signaled by the corresponding dynamic hysteresis loops on Fig. 3 of Ref. MRSh_theory_focused . We can conclude then that, for the same reason (lack of the linear response), the perpendicular orientation of DC and AC magnetic fields are in favor in the low-frequency limit both for the MRSh and sLLG cases.
Let us note that the low-frequency limit receives important application in the so-called image-guided thermal therapy by magnetic particle imaging (MPI); see, e.g. recent_focused_hyperthermia ; MRSh_theory_focused ; adv_func_mat ; review_mnp . MPI is a tomographic imaging method gleich that is based on the magnetization response of MNPs to generate an image. This provides high spatial and temporal resolution with excellent contrast (see, e.g., herz ; vaalma ) and its applicability to cancer imaging has also been demonstratedyu . The other advantage of the method is its combination with magnetic hyperthermia to provide real-time feedback and to overcome the nonspecific heating problem MRSh_theory_focused ; recent_focused_hyperthermia ; mpi_test . Basically, the same AC magnetic field can be used both for MPI and magnetic heating but with the requirement of low frequencies for MPI.
@@ II The MRSh equation
In cancer therapy by magnetic nanoparticle hyperthermia, MNPs are injected into the bloodstream during the process and allowed to collect at the tumor site because of the EPR effect. Once the MNPs are in position, they are exerted by an external AC magnetic field causing them to quickly heat up and destroy the cancer cells that encircle them. The application has limitations. The frequency cannot be chosen too high to minimize the eddy currents; a typical maximum value is around 5001000500-1000 kHz. In addition to that, the product of the frequency and the amplitude of the applied field, which is proportional to the power injected, has an upper bound, the Hergt-Dutz limit hergt_dutz . Thus, amplitudes are typically in the range of 103010-30 kA/m, while frequencies are in the range 100500100-500 kHz. Although the diameter of the MNP matters usov_claster , we do not consider this effect here and leave it for future work. In general, a good choice is between 10 and 50 nm which we use in our analysis. In addition to that, here we restrict the theoretical study to the isotropic case where no shape and no crystal anisotropy are considered for simplicity. Taking into account anisotropy would be a logical next step, but it is out of the scope of this paper.
As we argued, single domain MNPs dissipate heat due to Brown (rotation of the particle) or Néel (rotation of the magnetic moment) relaxation losses. To decide which process dominates, one has to consider the parameters: e.g. for high frequency ($f\gtrsim 100$ kHz) and small diameter ($\sim 20$ nm), only the orientation of their magnetic moment has to be taken into account usov_hysteresis ; lyutyy_general ; neel_brown ; viscous_rotating .
Calculation of the energy loss in a single cycle requires time-dependent magnetization $\bf m$}, which can be obtained by solving the phenomenological magnetization equation derived by Martsenyuk, Raikher, and Shliomis, i.e. the MRSh equation mrsh . This equation is valid at moderate to high field amplitudes and frequencies and is typically used to describe Brownian relaxation,
@@
where the dimensionless magnetization $M=m/m_{S}$ is used, and $H_{e}(t)$, called the effective magnetic field is calculated based on the assumption that the magnetization, $M(t)$ at time $t$, would be in equilibrium if the only external field present was $H_{e}(t)$.
In this work, Eq. (3) is solved to obtain the average of the magnetization vector component $M(t)$ that is parallel to the unidirectional alternating magnetic field.
The parameters are chosen to be $\eta_{B}=10^{-3}$ Pa s, d=20{d=20} nm, $\beta=2$ nm, $T=300$ K, $m_{S}=4.5\times 10^{5}$ A/m, for easy comparison with previous works such as MRSh_theory_focused where the same values are given, and are typical choices for a single crystal Fe3O4 Fannin . This yields the Brownian relaxation time $\tau_{B}\approx 5.25\times 10^{-6}$ s.
@@ III The stochastic LLG equation
Another option to obtain the time-dependent magnetization is the use of the sLLG which describes the motion of the magnetization vector of a single (isotropic) MNP that undergoes the Néel relaxation. The stochastic dynamics has been studied for a long time, supported by experimental works, e.g., thermal_exp and summarized in detail in reviews, e.g., Ref. thermal_summary . The sLLG equation sLLG is
$\frac{\rm{d}{\bf M}}{{\rm{d}}t}=-\gamma^{\prime}[{\bf M\times(H_{\rm{eff}}+H)}]+\alpha^{\prime}[[{\bf M\times(H_{\rm{eff}}+H)]\times M}]$ (4)
where thermal fluctuations are taken into account by the random magnetic field, ${\bf H}=(H_{x},H_{y},H_{z})$ where its components are independent Gaussian white noise variables,
@@
with $i,j=x,y,z$ and $\gamma^{\prime}=\mu_{0}\gamma_{0}/(1+\alpha^{2})$, $\alpha^{\prime}=\gamma^{\prime}\alpha$ with dimensionless damping $\alpha=\mu_{0}\gamma_{0}\eta_{N}$ with damping factor $\eta_{N}$ and $\gamma_{0}=1.76\times 10^{11}$ Am2/Js is the gyromagnetic ratio, while $\mu_{0}=4\pi\times 10^{-7}$ Tm/A (or N/A2) is the vacuum permeability. Here ${\bf M}={\bf m}/m_{S}$ is a 3-dimensional unit vector, which is normalized by its saturation magnetization, $m_{S}$. The parameter $D$ corresponds to the fluctuation-dissipation theorem, e.g., in Ref. path_int_sllg , and is defined as $D=\eta_{N}k_{B}T/(m_{S}V\mu_{0})$ with the Boltzmann factor, $k_{B}$, the absolute temperature, $T$, and the volume of the particle, $V$. The angular brackets stand for averaging over all possible realization of the stochastic field, ${\bf H}(t)$, and $\delta(t)$ is the Dirac $\delta$ function. This results in the Néel relaxation time
@@
In order to solve the sLLG equation numerically as an Ito process, it is useful to rewrite it in terms of spherical coordinates. This also ensures that the magnitude of the magnetization remains unchanged. In a simple case of a unidirectional alternating magnetic field the sLLG equation takes the form
@@
where the role of random white noise is played by the Gaussian random variables $n_{\phi}$ and $n_{\theta}$, which satisfy the relations
@@
Let us also introduce the frequency-like parameters, $\omega_{L}=H\gamma^{\prime}$ and $\alpha_{N}=H\alpha^{\prime}$, where the amplitude of the external magnetic field $H$ is also incorporated. The damping parameter can be chosen as $\alpha=0.1$ or $\alpha=0.3$, which are used in Ref. Lyutyy_energy and in Ref. Giordano , respectively. In order to enable a direct comparison with the results of MRSh, the value of $\alpha$ is changed in this work. If $\alpha=1.0$, the two models become qualitatively identical. The $m_{S}$ and $T$ are chosen to be the same as the ones listed in the previous section. However, it is important to note that, due to the different mechanisms, the relaxation times differ $\tau_{N}=1.5\times 10^{-7}\text{s}\neq\tau_{B}$}. To be in the Néel relaxation range, the volume of the magnetic particle must also be different, $V\approx 2.43\times 10^{-22}$ m3. With these parameters $\omega_{L}$ and $\alpha_{N}$ can be obtained.
@@ IV.1 Energy loss (SAR, ILP)
Let us note, that Eq. (9) contains the average magnetization, $\bf M$, of a single particle. The AC and DC magnetic fields are incorporated in the effective magnetic field $H_{\rm{eff}}$. It is often useful to introduce a dimensionless ratio,
@@
with $H$ being the amplitude of the AC magnetic field. By dividing by the field amplitude and the saturation magnetization, this quantity depends only on the shape of the dynamic hysteresis loop.
@@
ILP is used to compare SAR values obtained for different values of $f$ and $H$.
In this work $\phi=1$ is used for simplicity, while the density $\rho=5\times 10^{6}$ g/m3 is chosen, which is a typical value for single crystal Fe3O4 magnetite Ilg .

### v2

@@ I Introduction
Magnetic nanoparticle hyperthermia is an adjuvant treatment strategy in cancer therapy where injected magnetic nanoparticles (MNPs) are used to increase the temperature locally, see e.g., PRA1 ; PRA2 ; PMB1 ; PMB2 ; pankhurst ; ortega_d ; pankhurst_progress ; perigo_e_a ; cabrera_d ; clinapp_1 ; clinapp_2 ; review_mnp . The capacity of metallic nanoparticles to preferentially accumulate in tumor tissues is the primary benefit of using them in hyperthermia. This is due to the tumor’s enhanced permeability and retention (EPR) impact, which enables nanoparticles to penetrate and accumulate in the tumor location, for more details and references, see e.g., thesis_mehmeti . The abnormal design of tumor blood vessels, which are leaky and have irregular shapes, causes the EPR effect, enabling nanoparticles to travel through the gaps and collect in the tumor tissue. However, there are several difficulties and limitations linked to magnetic nanoparticle hyperthermia. One of the most difficult tasks is optimizing the nanoparticle characteristics and external energy source to obtain the best therapeutic results. This involves finding the best MNPs size, form, surface coating, and concentration, as well as the best energy dosage and frequency.
[…]
@@
Another challenge is to ensure the safety of the procedure, as overheating can harm healthy tissues and result in adverse effects such as discomfort, inflammation, or tissue necrosis. Thus, there is a need for spatially focused heating, which can be done by using the combination of static gradient (DC) and time-dependent alternating (AC) magnetic fields, where the temperature increase is observed only where the static field vanishes mpi_test . The improvement of the efficiency and the spatial focusing ability of MNP hyperthermia receives considerable attention, see e.g., adv_func_mat ; review_mnp and non_local_heat ; focused_hyperthermia ; focused_hyperthermia_2 ; focused_hyperthermia_3 ; focused_hyperthermia_4 ; recent_focused_hyperthermia ; stat_along_increased . It is known that the transferred energy is given by a bell-shaped curve as a function of the DC field amplitude, see e.g., Fig. (10.4) of review_mnp or superlocal_jmmm ; superlocal_jpd and for more details MRSh_theory_focused .
@@
The energy transferred in a single cycle of the AC field is identical to the area of the dynamic hysteresis loop, and it is related to the imaginary part of the frequency-dependent susceptibility. Measurements of dynamic hysteresis loops under the influence of AC and DC magnetic fields were reported in several publications, see for example Mehdaoui ; Myrovali ; Onodera . Similarly, a theoretical study of the dynamics of magnetization in the presence of AC and DC fields was carried out, for example in Dejardin ; Mrabti ; Murase ; MRSh_theory_focused , and recently, the study of the influence of static magnetic field on the specific absorption rate of the randomly oriented assembly of magnetic nanoparticles was reported in Rytov . In addition, the influence of anisotropy and magneto-dipole interaction effects was investigated in the presence of AC and DC fields, see e.g., usov_claster ; Zhao ; Usov_2019 ; viscous_rotating ; Engelmann ; hip_mc .
@@
The single-domain MNP dissipates heat as a result of relaxation losses under the applied AC magnetic field. This heat generation occurs due to either the physical rotation of the particle as a whole (Brownian relaxation) or the internal rotation of the magnetic dipole (Néel relaxation) while the particle is assumed to be fixed brown_neel_relax_1 ; brown_neel_relax_2 ; brown_neel_relax_3 . Although various mathematical approaches are necessary for a theoretical description of the energy absorption processes of immobilized and freely rotating nanoparticles, the usual theoretical framework is the Martsenyuk-Raikher-Shliomis (MRSh) mrsh equation for the Brownian, and the stochastic Landau-Lifshitz-Gilbert (sLLG) sLLG equation for the Néel relaxation.
@@
To do that extension, one possible choice is to use the MRSh equation and compare previous results of the sLLG and the MRSh equations directly. Our reasons for this choice are the following. It is well-known that under some constraints, the two different regimes, i.e. the Brownian and the Néel relaxations are very much separated by each other because the relaxation time $\tau$ is either dominated by $\tau_{{\mathrm{B}}}$ (Brown relaxation time) or by $\tau_{{\mathrm{N}}}$ (Néel relaxation time). So, it is a good approximation to use either the MRSh or the sLLG equations. However, by using the concept of magnetic and ordinary viscosity, the results of the sLLG equation can, in principle, be used to describe not just the motion of the magnetic dipole but also the particle. Of course, in this case the parameters (damping factor, frequency, magnetic field amplitude etc.) must be chosen appropriately, and these are not necessarily the same if one wants to obtain the same results by the two equations.
Indeed, one of our goals here is to show that one can reproduce well-known dynamic hysteresis loops and bell-shaped spatially-focused energy loss results of the MRSh approach by the sLLG equation. In particular, we demonstrate this by reproducing the MRSh results given on Fig. 2 and Fig. 3 of Ref. MRSh_theory_focused in the framework of the sLLG equation. For example, the spatial focusing ability is different in the right (low frequencies) and left (high frequencies) panels of Fig. 2 of Ref. MRSh_theory_focused . In this work, we show that this is the consequence of the use of a parallel combination of DC and AC magnetic fields and the breakdown of the linear response theory for low frequencies which is signaled by the corresponding dynamic hysteresis loops on Fig. 3 of Ref. MRSh_theory_focused . We can conclude then that for the same reason (lack of the linear response), the perpendicular orientation of DC and AC magnetic fields are in favor in the low-frequency limit both for the MRSh and sLLG cases.
Let us note that the low-frequency limit receives important application in the so-called image-guided thermal therapy by magnetic particle imaging (MPI), see e.g., recent_focused_hyperthermia ; MRSh_theory_focused ; adv_func_mat ; review_mnp . MPI is a tomographic imaging method gleich that is based on the magnetization response of MNPs to generate an image. This provides high spatial and temporal resolution with excellent contrast (see e.g., herz ; vaalma ) and its applicability to cancer imaging has also been demonstrated yu . The other advantage of the method is its combination with magnetic hyperthermia to provide real-time feedback and to overcome the nonspecific heating problem MRSh_theory_focused ; recent_focused_hyperthermia ; mpi_test . Basically, the same AC magnetic field can be used both for MPI and magnetic heating but with the requirement of low frequencies for MPI.
@@ II The MRSh equation
In cancer therapy by magnetic nanoparticle hyperthermia, MNPs are injected into the bloodstream during the process and allowed to collect at the tumor site because of the EPR effect. Once the MNPs are in position, they are exerted by an external AC magnetic field causing them to quickly heat up and destroy the cancer cells that encircle them. The application has limitations. The frequency cannot be chosen too high to minimize the eddy currents; a typical maximum value is around 5001000500-1000 kHz. In addition to that, the product of the frequency and the amplitude of the applied field, which is proportional to the power injected, has an upper bound, the Hergt-Dutz limit hergt_dutz ; hergt_dutz_2 ; hergt_dutz_3 . Thus, amplitudes are typically in the range of 104010-40 mT $\sim$ 103010-30 kA/m, while frequencies are in the range 100500100-500 kHz. Although the diameter of the MNP matters usov_claster , we do not consider this effect here and leave it for future work. In general, a good choice is between 10 and 50 nm which we use in our analysis. In addition to that, here we restrict the theoretical study to the isotropic case where no shape and no crystal anisotropy are considered for simplicity. Taking into account anisotropy would be a logical next step, but it is out of the scope of this paper.
As we argued, single domain MNPs dissipate heat due to Brown (rotation of the particle) or Néel (rotation of the magnetic moment) relaxation losses. To decide which process dominates, one has to consider the parameters: e.g., for high frequency ($f\gtrsim 100$ kHz) and small diameter ($\sim 20$ nm), only the orientation of their magnetic moment has to be taken into account usov_hysteresis ; lyutyy_general ; neel_brown ; viscous_rotating .
Calculation of the energy loss in a single cycle requires time-dependent magnetization $\bf m$}, which can be obtained by solving the phenomenological magnetization equation derived by Martsenyuk, Raikher, and Shliomis, i.e. the MRSh equation mrsh . This equation is valid at moderate high field amplitudes and frequencies and is typically used to describe Brownian relaxation,
@@
where the dimensionless magnetization $M=m/m_{S}$ is used, and $H_{e}(t)$ called the effective magnetic field is calculated based on the assumption that the magnetization, $M(t)$ at time $t$, would be in equilibrium if the only external field present was $H_{e}(t)$.
In this work Eq. (3) is solved to obtain the average of the magnetization vector component $M(t)$ that is parallel to the unidirectional alternating magnetic field.
The parameters are chosen to be: $\eta_{B}=10^{-3}$ Pa s, d=20{d=20} nm, $\beta=2$ nm, $T=300$ K, $m_{S}=4.5\times 10^{5}$ A/m, for an easy comparison with previous works such as MRSh_theory_focused , where the same values are given (typical choices for a single crystal Fe3O4 Fannin ). This yields the Brownian relaxation time $\tau_{B}\approx 5.25\times 10^{-6}$ s.
@@ III The stochastic LLG equation
Another option to obtain the time-dependent magnetization is the use of the sLLG, which describes the motion of the magnetization vector of a single (isotropic) MNP that undergoes the Néel relaxation. The stochastic dynamics has been studied for a long time, supported by experimental works, e.g., thermal_exp and summarized in detail in reviews, e.g., Ref. thermal_summary . The sLLG equation sLLG ; sLLG_2 ; sLLG_3 ; sLLG_4 ,
$\frac{\rm{d}{\bf M}}{{\rm{d}}t}=-\gamma^{\prime}[{\bf M\times(H_{\rm{eff}}+H)}]+\alpha^{\prime}[[{\bf M\times(H_{\rm{eff}}+H)]\times M}],$ (4)
where thermal fluctuations are taken into account by the random magnetic field ${\bf H}=(H_{x},H_{y},H_{z})$, where its components are independent Gaussian white noise variables,
@@
with $i,j=x,y,z$ and $\gamma^{\prime}=\mu_{0}\gamma_{0}/(1+\alpha^{2})$, $\alpha^{\prime}=\gamma^{\prime}\alpha$ with dimensionless damping $\alpha=\mu_{0}\gamma_{0}\eta_{N}$ with damping factor $\eta_{N}$ and $\gamma_{0}=1.76\times 10^{11}$ Am2/Js is the gyromagnetic ratio, while $\mu_{0}=4\pi\times 10^{-7}$ Tm/A (or N/A2) is the vacuum permeability. Here ${\bf M}={\bf m}/m_{S}$ is a 3-dimensional unit vector, which is normalized by its saturation magnetization, $m_{S}$. The parameter $D$ corresponds to the fluctuation-dissipation theorem, e.g., in Ref. path_int_sllg , and is defined as $D=\eta_{N}k_{B}T/(m_{S}V\mu_{0})$ with the Boltzmann factor $k_{B}$, the absolute temperature $T$, and the volume of the particle $V$. The angular brackets stand for averaging over all possible realization of the stochastic field, ${\bf H}(t)$, and $\delta(t)$ is the Dirac $\delta$ function. This results in the Néel relaxation time,
@@
In order to solve the sLLG equation numerically as an Ito process, it is useful to rewrite it in terms of spherical coordinates. This also ensures that the magnitude of the magnetization remains unchanged. In a simple case of a unidirectional alternating magnetic field the sLLG equation takes the form:
@@
where the role of random white noise is played by the Gaussian random variables $n_{\phi}$ and $n_{\theta}$, which satisfy the relations,
@@
Let us also introduce the frequency-like parameters, $\omega_{L}=H\gamma^{\prime}$ and $\alpha_{N}=H\alpha^{\prime}$, where the amplitude of the external magnetic field $H$ is also incorporated. The damping parameter can be chosen as $\alpha=0.1$ or $\alpha=0.3$, which are used in Ref. Lyutyy_energy and in Ref. Giordano respectively. In order to enable a direct comparison with the results of MRSh, the value of $\alpha$ is changed in this work. If $\alpha=1.0$, the two models become qualitatively identical. The $m_{S}$ and $T$ are chosen to be the same as the ones listed in the previous section. However, it is important to note that, due to the different mechanisms, the relaxation times differ $\tau_{N}=1.5\times 10^{-7}\text{s}\neq\tau_{B}$}. To be in the Néel relaxation range, the volume of the magnetic particle must also be different, $V\approx 2.43\times 10^{-22}$ m3. With these parameters $\omega_{L}$ and $\alpha_{N}$ can be obtained.
@@ IV.1 Energy loss (SAR, ILP)
Let us note, that Eq. (9) contains the average magnetization $\bf M$, of a single particle. The AC and DC magnetic fields are incorporated in the effective magnetic field $H_{\rm{eff}}$. It is often useful to introduce a dimensionless ratio,
@@
with $H$ being the amplitude of the AC magnetic field. Dividing by the field amplitude and the saturation magnetization, this quantity depends only on the shape of the dynamic hysteresis loop.
@@
ILP is used to compare SAR values obtained for different values of $f$ and $H$. In this work $\phi=1$ is used for simplicity, while the density $\rho=5\times 10^{6}$ g/m3 is chosen, which is a typical value for single crystal Fe3O4 magnetite Ilg .